OpenStudy (anonymous):

find out for what value of p the improper integral +∞ ∫ 1/x^p dx converge. 1 tough questions,,,, i need an explaintion :((

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

You need to consider specific cases for p. You set the problem up like this\[\lim_{c \rightarrow \infty} \int\limits_{1}^{c}\frac{1}{x^p}dx\]and consider when: 1. p=1 2. p<1 3. p>1 The integral will only converge for p<1, but I'll leave you to have a go at the mathematics. If you need anything else, let me know.

OpenStudy (anonymous):

We have to set the integral up as above since this is Riemannian integration, which is only defined on intervals not including +/- infinity. Once we integrate, that's when we can use limits to determine the value.

OpenStudy (anonymous):

THANKS ALOT

OpenStudy (anonymous):

Become a fan :)

OpenStudy (anonymous):

but that doesnt make sense, i just did the math, it converges only when p>1 ....